# DiffMesh: A Motion-aware Diffusion Framework for Human Mesh Recovery from Videos

### v4

@@ Abstract
Human mesh recovery (HMR) provides rich human body information for various real-world applications such as gaming, human-computer interaction, and virtual reality. While image-based HMR methods have achieved impressive results, they often struggle to recover humans in dynamic scenarios, leading to temporal inconsistencies and non-smooth 3D motion predictions due to the absence of human motion. In contrast, video-based approaches leverage temporal information to mitigate this issue. In this paper, we present DiffMesh, an innovative motion-aware diffusion framework for video-based HMR. DiffMesh establishes a bridge between diffusion models and human motion, efficiently generating accurate and smooth output mesh sequences by incorporating human motion within the forward process and reverse process in the diffusion model. Extensive experiments are conducted on the widely used datasets (Human3.6M [16] and 3DPW [51]), which demonstrate the effectiveness and efficiency of our DiffMesh. Visual comparisons in real-world scenarios further highlight DiffMesh’s suitability for practical applications. The code will be publicly available.
Figure 1: Different approaches of applying diffusion model for video-based HMR, the input frame $f$ is 3 for simplicity and the number of steps is $N$. Here $x_{i}$ and $c_{i}$ denote the mesh and conditional features of $i_{th}$ frame, respectively. (a) Baseline 1: The diffusion model is applied for each frame individually. The total steps are $f\times N$, which is computationally expensive. Moreover, motion is ignored during the denoising process, leading to non-smooth motion predictions (high acceleration error). (b) Baseline 2: The features from ground truth mesh vertices of each frame $x_{i}$ are concatenated to unified features $x_{unified}$ during the forward process. To obtain the mesh of each frame, features are split after the denoising process. Although this strategy reduces the total steps to $N$, it doesn’t effectively model human motion during denoising (acceleration error can be further reduced). (c) Our proposed DiffMesh: We consider the inherent motion patterns within the forward process and the reverse process when utilizing the diffusion model, leading to smooth and accurate motion predictions (low acceleration error and MPVE). The total number of steps remains as $N$. For more detailed information, please refer to Sec 3.3.
@@ 1 Introduction
While it is straightforward to apply the diffusion model for image-based HMR such as in [11], effective utilization of the diffusion model for video-based HMR remains a formidable challenge. As previous video-based HMR methods have not yet incorporated diffusion models into their frameworks, we provide a summary of diffusion-based methods related to human pose and mesh recovery (please refer to Sec. 2.2 for further elaboration). We then organize these methods into two intuitive baseline approaches that can be applied for video-based HMR, as illustrated in Fig. 1 (a) and (b). Assuming a video input with $f$ frames in Baseline 1, the diffusion model is applied individually to recover the mesh of the single frame (usually would be the center frame). More details can be found in Sec. 3.2. However, this approach suffers from significant computational burdens as the total number of denoising steps amounts to $f\times N$ to recover the mesh sequence of $f$ frames, with $N$ representing the number of denoising steps in the diffusion model. Moreover, the diffusion model does not account for motion smoothness in dynamic sequences, which may result in potentially temporal inconsistent mesh predictions. To mitigate computational costs and enhance inference speed, we introduce Baseline 2. The features for ground truth mesh vertices of each frame $x_{i}$, where $i\in\left\{1,\dots f\right\}$ are concatenated to unified features $x_{unified}$ before the forward diffusion process. Subsequently, the output features are partitioned back into individual frames $x_{i}$ after the denoising process. While this strategy effectively reduces the steps from $f\times N$ to $N$, it is important to note that motion smoothness in dynamics is still not considered during the forward and reverse process since features across frames are simply concatenated. This limitation can potentially lead to non-smooth motion predictions.
@@ 2.1 Video-based HMR
Image-based HMR methods  [32, 68, 25, 66, 62, 30, 12, 53, 19] have achieved remarkable progress in accuracy. However, when applied to video sequences, these image-based methods suffer from severe motion jitter due to frame-by-frame reconstruction, making them unsuitable for practical use. To address this issue, video-based methods [18, 34, 47, 60, 26, 60, 45, 59, 56] utilize temporal information to enhance accurate and temporally consistent human mesh from video frames. VIBE [22] leverages the AMASS dataset [35] to discriminate between real human motions and those produced by its temporal encoder. MEVA [34] estimates the coarse 3D human motion using a variational motion estimator, then refines the motion by a motion residual regressor. TCMR [5] utilizes GRU-based temporal encoders to forecast the current frame from the past and future frames, which enhances the temporal consistency. Similar to [5], MPS-Net [54] captures the motion continuity dependencies by the attention mechanism. MotionBERT [70] proposes a pertaining stage to recover 3D motion from noise 2D observations, achieving state-of-the-art performance.
@@ 2.2 Diffusion Generative Models
Diffusion Generative Models have achieved impressive success in a wide variety of computer vision tasks such as image inpainting [46], text-to-image generation [43], and image-to-image translation [7]. Given the strong capability to bridge the large gap between highly uncertain and determinate distribution, several works have utilized the diffusion generative model for the text-to-motion generation [65, 48, 64], human pose estimation [44], object detection [4], and head avatar generation [2, 37]. DiffPose [13] is the first to utilize a diffusion-based approach for predicting the 3D pose from a 2D pose sequence input. Shan et al. [44] generate multiple hypotheses of the 3D pose from 2D pose input using a diffusion model, and then a multi-hypothesis aggregation module outputs the final 3D pose. EgoEgo [27] employs a diffusion model to generate multiple plausible full-body motions based on the head pose input.
@@ 3.2 Diffusion-based Baselines
Baseline 1: As illustrated in Fig. 4 (a), we initially apply the vanilla diffusion model at the frame level. For each $i_{th}$ frame, we employ a backbone to extract the conditional feature $c_{i}$ where $i\in\left\{1,\dots,f\right\}$ and $f$ representing the total number of frames. We then proceed with the standard forward diffusion process, introducing noise to the ground truth mesh vertices $x_{i}$. After $N$ steps, the feature $x_{i}^{end}$ converges to a Gaussian Noise. Subsequently, a transformer-based diffusion model (details are included in the supplementary Sec. 2) is employed for the denoising process. The denoised feature $x^{\prime}_{i}$ is obtained after $N$ denoising steps. Finally, the human mesh of $i_{th}$ frame is returned by the human mesh head. By iterating this procedure $f$ times, we generate the final human mesh sequence consisting of $f$ frames. However, despite its simplicity and straightforwardness, Baseline 1 exhibits a significant drawback: the total number of steps is $f\times N$, resulting in considerable computational overhead and slow inference speed. Furthermore, it fails to address motion smoothness within the diffusion model, potentially leading to temporal inconsistencies in mesh predictions.
Baseline 2: To mitigate computational costs and enhance inference speed, we introduce Baseline 2 for video-based HMR. Motivated by recent works [65, 20, 27] that leverage diffusion models for motion generation, we adopt a similar concatenation and separation strategy, as illustrated in Fig. 4 (b). In this approach, we concatenate the ground-truth mesh vertices of each frame into a unified feature $x_{unified}=cat(x_{1},x_{2},\dots,x_{f})$, where $f$ represents the total number of frames. Similarly, we create a concatenated conditional feature $c_{unified}$ by combining the image features from each frame. Subsequently, we apply the forward diffusion process by introducing noise for $N$ steps, which is analogous to Baseline 1. After $N$ steps, $x_{unified}^{end}$ also converges to a Gaussian Noise state. Following this, we perform $N$ denoising steps utilizing a transformer-based diffusion model (details are included in the supplementary Sec. 2). The denoised feature $x^{\prime}_{unified}$ is then partitioned into $x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{f}$. Finally, the mesh sequence is obtained through the mesh head.
@@ 3.3 DiffMesh
Forward Process: We make the assumption that human motion consistently influences the human mesh vertices (in a latent space) across frames, denoted as $x_{i}$ to $x_{i+1}$, where $i\in\left\{1,\dots f-1\right\}$. This motion pattern can be conceptualized as a specific noise introduced to the mesh vertices feature of the preceding frame:
@@
where $m_{i}$ represents the motion pattern from $i_{th}$ frame to $(i+1)_{th}$ frame. Consequently, $x_{1}$ can be gradually diffused into $x_{f}$ after $f-1$ forward diffusion steps. Importantly, achieving $x_{f}$ does not denote the conclusion of the forward diffusion process. We assume that human motion will eventually reach the initial distribution $x_{end}$ through total $N$ diffusion steps (similar to the total $N$ steps to get the Gaussian distribution in the conventional diffusion model). More details and ablation study are provided in the supplementary Sec. 3. To arrive at this initial distribution $x_{end}$, DiffMesh requires $N-f+1$ additional forward diffusion steps from $x_{f}$. Therefore, the total number of steps remains $N$.
@@
Given the input feature $x^{\prime}_{i}$ and corresponding conditional feature $c_{i}$ at each step, our network produces the predicted motion $m_{i-1}$ as well as the predicted previous conditional feature $\hat{c}_{i-1}$. The feature of initial distribution (mesh template) is denoted as $x_{end}$. Similar to Equation 5, the mesh features $x^{\prime}_{i-1}$ in DiffMesh can be computed as
$\small x^{\prime}_{i-1}=\frac{1}{\sqrt{\beta_{i}}}(x^{\prime}_{i}-\frac{1-\beta_{i}}{\sqrt{1-\hat{\beta_{i}}}}{\color[rgb]{1,.5,0}m_{i-1}})+\sigma_{t}{\color[rgb]{1,.5,0}x_{end}},$ (7)
@@ 3.4 Overall Training and Loss Functions
Based on the assumption that human motion consistently impacts the human mesh across consecutive frames, we postulate the existence of a latent space where the motion pattern $m_{i}$ can be conceptualized as a specific noise introduced to the mesh vertices’ features of the preceding frame $x_{i}$, as depicted in Eq. 6. Our objective is to derive such motion patterns for each time interval within this specific latent space during the reverse steps. Ultimately, the final mesh sequence is generated based on the estimated mesh features $x_{i}$ through MLP layers and the SMPL model.
Thus, besides supervising the diffusion noise in the diffusion models, we also supervise the SMPL parameters of each frame ($\Theta_{i}$ and $\beta_{i}$) where $i\in\left\{1,\dots f\right\}$ with the ground-truth SMPL parameters using the $L2$ loss during training. Following [22, 5], the 3D joint coordinates are regressed by forwarding the SMPL parameters to the SMPL human body model. Furthermore, as mentioned in Sec. 3.3, Mean Squared Error (MSE) loss is applied between the conditional feature $c_{i}$ and the predicted conditional features $\hat{c_{i}}$. More detailed information is included in the supplementary Sec. 2.
@@ 5 Conclusion
In this paper, we present a novel diffusion framework (DiffMesh) for human mesh recovery from a video. DiffMesh innovatively connects diffusion models with human motion, resulting in the efficient generation of highly precise and seamlessly smooth output mesh sequences. Extensive experiments are conducted on Human3.6M and 3DPW datasets showing impressive performance compared to the SOTA video-based methods.
@@ Appendix A Overview
  • Section C: More Experiments about the number of input frames, the two-stream transformer network, initial distributions, and the auxiliary loss.
@@ B.3 Loss Function
This auxiliary loss contributes to the refinement of our transformer-based diffusion model during the training process. Thus, the overall loss for our DiffMesh is the sum of the $\mathcal{L}_{HMR}$, $\mathcal{L}_{aux}$ and $\mathcal{L}_{diffusion}$:
@@ B.4 More Details about the Architecture
Appendix C More Experiments
C.1 Performance on MPI-INF-3DHP dataset

### v6

@@ Abstract
Human mesh recovery (HMR) provides rich human body information for various real-world applications such as gaming, human-computer interaction, and virtual reality. While image-based HMR methods have achieved impressive results, they often struggle to recover humans in dynamic scenarios, leading to temporal inconsistencies and non-smooth 3D motion predictions due to the absence of human motion. In contrast, video-based approaches leverage temporal information to mitigate this issue. In this paper, we present DiffMesh, an innovative motion-aware diffusion framework for video-based HMR. DiffMesh establishes a bridge between diffusion models and human motion, efficiently generating accurate and smooth output mesh sequences by incorporating human motion within the forward process and reverse process in the diffusion model. Extensive experiments are conducted on the widely used datasets (Human3.6M [16] and 3DPW [53]), which demonstrate the effectiveness and efficiency of our DiffMesh. Visual comparisons in real-world scenarios further highlight DiffMesh’s suitability for practical applications. The project webpage is: https://zczcwh.github.io/diffmesh_page/
Figure 1: Different approaches of applying diffusion model for video-based HMR, the input frame $f$ is 3 for simplicity and the number of steps is $N$. Here $x_{i}$ and $c_{i}$ denote the mesh and conditional features of $i_{th}$ frame, respectively. (a) Baseline 1: The diffusion model is applied for each frame individually. The total steps are $f\times N$, which is computationally expensive. Moreover, motion is ignored during the denoising process, leading to non-smooth motion predictions. (b) Baseline 2: The features from ground truth mesh feature of each frame $x_{i}$ are concatenated to unified features $x_{unified}$ during the forward process. To obtain the mesh of each frame, features are split after the denoising process. Although this strategy reduces the total steps to $N$, it doesn’t effectively model human motion during denoising. (c) Our proposed DiffMesh: We consider the inherent motion patterns within the forward process and the reverse process when utilizing the diffusion model, leading to smooth and accurate motion predictions. The total number of steps remains as $N$. For more detailed information, please refer to Sec 3.3.
@@ 1 Introduction
While it is straightforward to apply the diffusion model for image-based HMR such as in [11], effective utilization of the diffusion model for video-based HMR remains a formidable challenge. As previous video-based HMR methods have not yet incorporated diffusion models into their frameworks, we provide a summary of diffusion-based methods related to human pose and mesh recovery (please refer to Sec. 2.2 for further elaboration). We then organize these methods into two intuitive baseline approaches that can be applied for video-based HMR, as illustrated in Fig. 1 (a) and (b). Assuming a video input with $f$ frames in Baseline 1, the diffusion model is applied individually to recover the mesh of the single frame (usually would be the center frame). More details can be found in Sec. 3.2. However, this approach suffers from significant computational burdens as the total number of denoising steps amounts to $f\times N$ to recover the mesh sequence of $f$ frames, with $N$ representing the number of denoising steps in the diffusion model. Moreover, the diffusion model does not account for motion smoothness in dynamic sequences, which may result in potentially temporal inconsistent mesh predictions. To mitigate computational costs and enhance inference speed, we introduce Baseline 2. The mesh feature of each frame $x_{i}$, where $i\in\left\{1,\dots f\right\}$ are concatenated to unified features $x_{unified}$ before the forward diffusion process. Subsequently, the output features are partitioned back into individual frames $x_{i}$ after the denoising process. While this strategy effectively reduces the steps from $f\times N$ to $N$, it is important to note that motion smoothness in dynamics is still not considered during the forward and reverse process since features across frames are simply concatenated. This limitation can potentially lead to non-smooth motion predictions.
@@ 2.1 Video-based HMR
Image-based HMR methods  [33, 70, 25, 68, 64, 31, 12, 55, 19] have achieved remarkable progress in accuracy. However, when applied to video sequences, these image-based methods suffer from severe motion jitter due to frame-by-frame reconstruction, making them unsuitable for practical use. To address this issue, video-based methods [18, 35, 48, 62, 26, 46, 61, 58, 72, 50] utilize temporal information to enhance accurate and temporally consistent human mesh from video frames. VIBE [22] leverages the AMASS dataset [36] to discriminate between real human motions and those produced by its temporal encoder. MEVA [35] estimates the coarse 3D human motion using a variational motion estimator, then refines the motion by a motion residual regressor. TCMR [5] utilizes GRU-based temporal encoders to forecast the current frame from the past and future frames, which enhances the temporal consistency. Similar to [5], MPS-Net [56] captures the motion continuity dependencies by the attention mechanism. MotionBERT [72] proposes a pertaining stage to recover 3D motion from noise 2D observations, achieving state-of-the-art performance.
@@ 2.2 Diffusion Generative Models
Diffusion Generative Models have achieved impressive success in a wide variety of computer vision tasks such as image inpainting [47], text-to-image generation [44], and image-to-image translation [7]. Given the strong capability to bridge the large gap between highly uncertain and determinate distribution, several works have utilized the diffusion generative model for the text-to-motion generation [67, 49, 66, 30], human pose estimation [45], object detection [4], and head avatar generation [2, 38]. DiffPose [13] is the first to utilize a diffusion-based approach for predicting the 3D pose from a 2D pose sequence input. Shan et al. [45] generate multiple hypotheses of the 3D pose from 2D pose input using a diffusion model, and then a multi-hypothesis aggregation module outputs the final 3D pose. EgoEgo [27] employs a diffusion model to generate multiple plausible full-body motions based on the head pose input.
@@ 3.2 Diffusion-based Baselines
Baseline 1: As illustrated in Fig. 4 (a), we initially apply the vanilla diffusion model at the frame level. For each $i_{th}$ frame, we employ a backbone to extract the conditional feature $c_{i}$ where $i\in\left\{1,\dots,f\right\}$ and $f$ representing the total number of frames. We then proceed with the standard forward diffusion process, introducing noise to the mesh feature $x_{i}$ from ground truth SMPL parameters. After $N$ steps, the feature $x_{i}^{end}$ converges to a Gaussian Noise. Subsequently, a transformer-based diffusion model (details are included in the supplementary Sec. 2) is employed for the denoising process. The denoised feature $x^{\prime}_{i}$ is obtained after $N$ denoising steps. Finally, the human mesh of $i_{th}$ frame is returned by the human mesh head. By iterating this procedure $f$ times, we generate the final human mesh sequence consisting of $f$ frames. However, despite its simplicity and straightforwardness, Baseline 1 exhibits a significant drawback: the total number of steps is $f\times N$, resulting in considerable computational overhead and slow inference speed. Furthermore, it fails to address motion smoothness within the diffusion model, potentially leading to temporal inconsistencies in mesh predictions.
Baseline 2: To mitigate computational costs and enhance inference speed, we introduce Baseline 2 for video-based HMR. Motivated by recent works [67, 20, 27] that leverage diffusion models for motion generation, we adopt a similar concatenation and separation strategy, as illustrated in Fig. 4 (b). In this approach, we concatenate the ground-truth mesh feature of each frame into a unified feature $x_{unified}=cat(x_{1},x_{2},\dots,x_{f})$, where $f$ represents the total number of frames. Similarly, we create a concatenated conditional features $c_{unified}$ by combining the image features from each frame. Subsequently, we apply the forward diffusion process by introducing noise for $N$ steps, which is analogous to Baseline 1. After $N$ steps, $x_{unified}^{end}$ also converges to a Gaussian Noise state. Following this, we perform $N$ denoising steps utilizing a transformer-based diffusion model (details are included in the supplementary Sec. 2). The denoised feature $x^{\prime}_{unified}$ is then partitioned into $x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{f}$. Finally, the mesh sequence is obtained through the mesh head.
@@ 3.3 DiffMesh
Forward Process: We make the assumption that human motion consistently influences the human mesh feature (in a latent space) across adjacent frames, denoted as $x_{i}$ to $x_{i+1}$, where $i\in\left\{1,\dots f-1\right\}$. This motion pattern can be conceptualized as a specific noise introduced to the mesh feature of the preceding frame:
@@
where $m_{i}$ represents the motion pattern from $i_{th}$ frame to $(i+1)_{th}$ frame. Consequently, the mesh feature of the first frame $x_{1}$ can be gradually diffused into the mesh feature of the last frame $x_{f}$ after $(f-1)$ forward diffusion steps. Importantly, achieving the last frame $x_{f}$ does not denote the conclusion of the forward diffusion process. We assume that mesh feature with adding human motion will eventually reach the initial distribution $x_{end}$ through total $N$ diffusion steps (similar to the total $N$ steps to get the Gaussian distribution in the conventional diffusion model). More details such as mathematical proof and ablation study are provided in the supplementary Sec. 3. To arrive at this initial distribution $x_{end}$, DiffMesh requires $N-f+1$ additional forward diffusion steps from $x_{f}$. Therefore, the total number of steps remains $N$.
@@
Given the input feature $x^{\prime}_{i}$ and corresponding conditional feature $c_{i}$ at each step, our network produces the predicted motion $\hat{m}_{i-1}$ as well as the predicted previous conditional feature $\hat{c}_{i-1}$. The feature of initial distribution is denoted as $x_{end}$. Similar to Equation 5, the mesh feature $x^{\prime}_{i-1}$ in DiffMesh can be computed as
$\small x^{\prime}_{i-1}=\frac{1}{\sqrt{\beta_{i}}}(x^{\prime}_{i}-\frac{1-\beta_{i}}{\sqrt{1-\hat{\beta_{i}}}}{\color[rgb]{1,.5,0}\hat{m}_{i-1}})+\sigma_{t}{\color[rgb]{1,.5,0}x_{end}},$ (7)
@@ 3.4 Overall Training and Loss Functions
Based on the assumption that human motion consistently impacts the human mesh across consecutive frames, we postulate the existence of a latent space where the motion pattern $m_{i}$ can be conceptualized as a specific noise introduced to the mesh feature of the preceding frame $x_{i}$, as depicted in Eq. 6. Our objective is to derive such motion patterns for each time interval within this specific latent space during the reverse steps. Ultimately, the final mesh sequence is generated based on the estimated mesh feature $x_{i}$ through MLP layers and the SMPL model.
We supervise the estimated SMPL parameters of each frame ($\Theta_{i}$ and $\beta_{i}$) where $i\in\left\{1,\dots f\right\}$ with the ground-truth using the $L2$ loss during training. Following [22, 5], the 3D joint coordinates are regressed by forwarding the SMPL parameters to the SMPL human body model. Furthermore, as mentioned in Sec. 3.3, Mean Squared Error (MSE) loss is applied between the conditional feature $c_{i}$ and the predicted conditional features $\hat{c_{i}}$. More detailed information is included in the supplementary Sec. 2.
@@ 5 Conclusion
In this paper, we present a novel diffusion framework (DiffMesh) for human mesh recovery from a video. DiffMesh innovatively connects diffusion models with human motion, resulting in the efficient generation of highly precise and seamlessly smooth output mesh sequences. Extensive experiments are conducted on Human3.6M and 3DPW datasets showing impressive performance.
@@ Appendix A Overview
  • Section C: Mathematical proof and more experiments about the number of input frames, the two-stream transformer network, initial distributions, and the auxiliary loss.
@@ B.3 Loss Function
This auxiliary loss contributes to the refinement of our transformer-based diffusion model during the training process. Thus, the overall loss for our DiffMesh is the sum of the $\mathcal{L}_{HMR}$ and $\mathcal{L}_{aux}$:
@@ B.4 More Details about the Architecture
Appendix C Mathematical Proof and More Experiments
C.1 Mathematical Proof of modeling human motion as noise in diffusion model
Our approach draws an analogy between human motion and noise, treating the motion between adjacent frames as a structured form of noise. By operating in a high-dimensional latent space, we capture the complexity of human motion, where small perturbations (or ”noise”) in this space can be modeled as Gaussian. This allows us to align the problem with the core principles of diffusion models.
Following Equation 6 in the main paper, we have
$\mathbb{E}_{x_{0}\sim q}\left[-\log p(x_{0})\right]=\mathbb{E}_{x_{0}\sim q}\left[\log\mathbb{E}_{x_{1:T},y_{0:T}\sim q}\frac{q(x_{1:T},y_{0:T}|x_{0})}{p(x_{0:T},y_{0:T})}\right]$
Using Jensen’s Inequality, we can bound Eq 11 by moving the expectation inside the logarithm:
$\mathbb{E}_{x_{0}\sim q}\left[-\log p(x_{0})\right]\leq\mathbb{E}_{x_{0},x_{1:T},y_{0:T}\sim q}\left[\log\frac{q(x_{1:T},y_{0:T}|x_{0})}{p(x_{0:T},y_{0:T})}\right]$
The forward process in DiffMesh is designed to model the human motion between adjacent frames as structured noise. This noise is Gaussian in the latent space, where the human motion patterns are simplified. We define the forward process for the latent motion as:
$q(m_{t+1}|m_{t})=\mathcal{N}(m_{t+1};\sqrt{1-\beta_{t}}m_{t},\beta_{t}I)$ (13)
Here $m_{t}$ represents the motion noise at time step $t$, and $\beta_{t}$ controls the level of noise added to the motion between consecutive frames. This equation models the motion between frames as Gaussian perturbations in latent space.
We now need to decompose the joint probabilities $q(x_{1:T},y_{0:T}|x_{0})$ and $p(x_{0:T},y_{0:T})$ into their respective transition probabilities over time $t$:
The forward process $q(x_{1:T},y_{0:T}|x_{0})$ can be written as:
$q(x_{1:T},y_{0:T}|x_{0})=q(x_{T}|x_{0})\prod_{t=2}^{T}q(x_{t-1}|x_{t},x_{0})\prod_{t=0}^{T}q(y_{t}|x_{t})$
Similarly, the reverse process $p(x_{0:T},y_{0:T})$ is:
$p(x_{0:T},y_{0:T})=p(x_{T})\prod_{t=T}^{1}p(x_{t-1}|x_{t})\prod_{t=T}^{0}p(y_{t}|x_{t})$
Next, we focus on deriving the bound based on KL divergence. Substituting the decomposed expressions, we obtain two main terms: one for the states $x$ and another for the observations $y$
$\mathbb{E}_{x_{0},x_{1:T},y_{0:T}\sim q}\left[\log\frac{q(x_{T}|x_{0})\prod_{t=2}^{T}q(x_{t-1}|x_{t},x_{0})}{p(x_{T})\prod_{t=T}^{1}p(x_{t-1}|x_{t})}\right]$
$+\mathbb{E}_{x_{0},x_{1:T},y_{0:T}\sim q}\left[\log\frac{\prod_{t=0}^{T}q(y_{t}|x_{t})}{\prod_{t=T}^{0}p(y_{t}|x_{t})}\right]$ (16)
Each of these terms corresponds to the forward process (adding Gaussian motion noise) and the reverse process (denoising to recover the original human motion).
The final step is to express the result as a sum of KL divergences. For the forward and reverse processes of both the states $x$ and the observations $y$, we can represent this as:
$=D_{\text{KL}}(q(x_{T}|x_{0})||p(x_{T}))+\mathbb{E}_{q}\left[-\log p(x_{0}|x_{1})\right]$
$+\sum_{t=2}^{T}D_{\text{KL}}(q(x_{t-1}|x_{t},x_{0})||p(x_{t-1}|x_{t}))$
$+\sum_{t=0}^{T}D_{\text{KL}}(q(y_{t}|x_{t})||p(y_{t}|x_{t}))$ (17)
Here $D_{\text{KL}}(q(x_{T}|x_{0})||p(x_{T}))$ measures the divergence between the final denoised human mesh and the actual motion, and $D_{\text{KL}}(q(x_{t-1}|x_{t},x_{0})||p(x_{t-1}|x_{t}))$ measures the divergence at each intermediate step in recovering the motion, helping to maintain temporal consistency. Thus, we can derive Equation 7 in the main paper.
C.2 Performance on MPI-INF-3DHP dataset